Question 3 (20 marks)

  1. Milton Electronics Inc. purchased inventory costing $48,000 by signing an 8-month, 6% note payable on Sept 1 2019. The note will be repaid with interest at maturity. Prepare journal entries to record the purchase of the inventory, accrual of interest on Dec 31, 2019 and the final repayment of the note at maturity. (10 marks)
  1. Milton Electronics Inc. started its business in 2019 and sold electronics worth $185,000 in 2019. The company provided a 2- year limited warranty for all sales. As per company estimate warranty costs would be 1% of sales in the first year and 3% of sales in the second year. By end of 2019 Milton Electronics Inc had already spent $1,510 on warranty repairs. Prepare all journal entries related to the warranty for 2019. Also Calculate the amount of Estimated Warranty Payable to be reported on the Balance Sheet on Dec 31, 2019 (10 marks)
Answer #1

a)

Sep1,19

Inventory $48000

To 6%notes payable $48000

(Being inventory purchased on credit)

31dec,19

Interest $960

To interest payable $ 960

(Being interest accrue but not paid)

30april,20

6% Notes payable. $48000 Interest payable* . $ 960    Interest* . $ 960

TO Cash A/C $49920

(Being notes payable with interest on maturity)

Notes:-

* Interest payable = notes payable* rate p.a.*no. Of months amount due

=48000* 6%*4/12

=$960

* Interest= notes payable*rate p.a.*no. Of months amount paid

=48000*6%*4/12

=$960

b)

Warranty expense $1850

To provision for warranty expense * $1850

(Being provision for 1% warranty expenses payable made)

Provision for Warranty Expenses $1510

To cash a/c $1510

(Being amount of warranty spent)

31dec,19

Estimated warranty liability* $ 5550

To provision for warranty expense $5550

(Being estimated liability recorded)

Notes:-

* 1% warranty payable= net sales*1%

=185000*1%

=$1850

*Estimated warranty liability= net sales*3%

=185000*3%

=$5550